Matt travels at a speed of 64km/hr, if Matt travelled at 78 km/hr he would arrive at his destination 10 Min earlier. What distance did Matt travel?
Answer by stanbon(75887)

Matt travels at a speed of 64km/hr, if Matt travelled at 78 km/hr he would arrive at his destination 10 Min earlier. What distance did Matt travel?
---------
Slower speed DATA:
rate = 64 km/hr ; distance = x km ; time = x/64 hrs
------------
Faster speed DATA:
rate = 78 km/hr ; distance = x km ; time = x/78 hrs
----------
Equation:
time - time = 1/6 hr
-------
x/64 - x/78 = 1/6
--------
6*78x - 6*64x = 64*78
86x = 64*78
----
x = 59.43 km
=================
Cheers,
Stan H.